# HoTPP Benchmark: Are We Good at the Long Horizon Events Forecasting?

### v2

@@ C.1 Architecture
HoTPP incorporates best practices from extensible and reproducible ML pipelines. It leverages PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) as the core training library, ensuring reproducibility and portability across various computing environments. Additionally, HoTPP utilizes the Hydra configuration library (Yadan, 2019) to enhance extensibility. The overall architecture is illustrated in Figure 8. HoTPP supports both discrete-time and continuous-time models, as well as RNN and Transformer architectures. Implementing a new method requires only the addition of essential components, while the rest can be configured through Hydra and configuration files.
Figure 8: The architecture of the HoTPP library.
C.2 Metrics
@@
MTPP models are typically evaluated based on the accuracy of next-event predictions. Common metrics include mean absolute error (MAE) or root mean square error (RMSE) for time shifts, and accuracy or error rate for label predictions. Some studies also assess test set likelihood as predicted by the model, but we do not include this measure as it is intractable for IFTPP and Next-K models. Previous works have advanced long-horizon evaluation using optimal transport distance (OTD) (Mei et al., 2019; Xue et al., 2023). In addition to these metrics, we evaluate the novel T-mAP metric, which addresses the shortcomings of previous metrics, as discussed in Section 3.
C.3 Backbones
@@
We use three types of architectures in our experiments. The first is the GRU network (Cho et al., 2014), which is one of the top-performing neural architectures for event sequences (Babaev et al., 2022). We also implement the continuous-time LSTM (CT-LSTM) from the NHP method (Mei & Eisner, 2017). Since CT-LSTM requires a specialized loss function and increases training time, we use it exclusively with the NHP method, preferring GRU in other cases. An additional advantage of GRU is that its output is equal to its hidden state, simplifying the estimation of intermediate hidden states for autoregressive inference starting from the middle of a sequence. Lastly, we implement the AttNHP continuous-time transformer model (Yang et al., 2022).
C.4 Datasets
For the first time, we combine domains such as financial transactions, social networks, and medical records into a single evaluation benchmark. Specifically, we provide evaluation results on a transactional dataset (Babaev et al., 2022), the MIMIC-IV medical dataset (Johnson et al., 2023), and social network datasets (Retweet (Zhao et al., 2015), StackOverflow (Jure, 2014), and Amazon (Jianmo, 2018)). These datasets represent diverse underlying processes: social network data is influenced by cascades (Zhao et al., 2015), medical records exhibit repetitive patterns, and transactional data reflects daily activities, combining regularity with significant uncertainty.
The dataset statistics are presented in Table 2. The Transactions dataset has the longest average sequence length and the largest number of classes, while MIMIC-IV contains the highest number of sequences. Retweet is of medium size, whereas Amazon and StackOverflow can be considered small datasets.
Table 2: Datasets statistics
Dataset Sequences Events Mean Length Mean Horizon Length Mean Duration Classes Transactions 50k 43.7M 875 9.0 719 203 MIMIC-IV 120k 2.4M 19.7 6.6 503 34 Retweet 23k 1.3M 56.4 14.7 1805 3 Amazon 9k 403K 43.6 14.8 22.1 16 StackOverflow 2k 138K 64.2 12.0 55.3 22
@@ C.4 Datasets
Transactions [^3] https://huggingface.co/datasets/dllllb/age-group-prediction, Retweet [^4] https://huggingface.co/datasets/easytpp/retweet, Amazon [^5] https://huggingface.co/datasets/easytpp/amazon, and StackOverflow [^6] https://huggingface.co/datasets/easytpp/stackoverflow datasets were obtained from the HuggingFace repository. Transactions data was released in competition and come with a free license [^7] https://www.kaggle.com/competitions/clients-age-group/data. Retweet, Amazon, and StackOverflow come with an Apache 2.0 license. MIMIC-IV is subject to PhysioNet Credentialed Health Data License 1.5.0, which requires ethical use of this dataset. Because of a complex data structure, we implement a custom preprocessing pipeline for the MIMIC-IV dataset.
@@
C.5 Performance improvements
The HoTPP benchmark provides highly optimized training and inference procedures for efficient evaluation on datasets containing up to tens of millions of events. First, we implement parallel RNN inference, which reuses computations when inference is initiated from multiple starting points within a sequence. Additionally, we optimize the code and apply PyTorch Script to the CT-LSTM model from NHP and the ODE model. CT-LSTM from EasyTPP and the ODE model from the original repository serve as baselines. The timing results are presented in Table 3. Experiments were conducted using a synthetic batch with a size of 64, a sequence length of 100, and an embedding dimension of 64. Evaluation was performed on a single Nvidia RTX 4060 GPU. The results show that HoTPP is 17 times faster at RNN autoregressive inference compared to simple prefix extension. HoTPP also accelerates CT-LSTM and ODE by 4 to 8 times. These optimizations significantly extend the applicability of the implemented methods to larger-scale datasets [^8] Experiments are documented in the “notebooks” folder of the HoTPP repository. Additionally, HoTPP offers a GPU implementation of the Hungarian algorithm, which also finds applications in computer vision.
Table 3: HoTPP computation speed improvements in terms of seconds per batch.
@@ Appendix D Training details
We trained each model using the Adam optimizer with a learning rate of 0.001 and a scheduler that reduces the learning rate by 20% after every 5 epochs. Gradient clipping was applied, with a maximum L2-norm set to 1.
We performed computations on NVIDIA V100 and A100 GPUs, with some smaller experiments conducted on an Nvidia RTX 4060. Each method was trained on a single GPU. The training time varied depending on the dataset and method, ranging from 5 minutes for RMTPP on the StackOverflow dataset to 40 minutes for the same method on the Transactions dataset, and up to 15 hours for NHP on Transactions. Multi-seed evaluations took approximately five times longer to complete.
@@ Appendix E Hyperparameters
We analyzed the behavior of OTD and T-mAP metrics depending on their parameters. The results are shown in Figure 9. The OTD cost has little effect on the ranking of methods but significantly influenced the metric scale. The choice of the $\delta$ parameter in T-mAP can, in some cases, affect method rankings, but as long as $\delta$ is not set too small or too large, the metric demonstrates stable rankings. We set the $\delta$ parameter to approximately 10-30% of the horizon duration.
Evaluation hyperparameters are presented in Table 4. Dataset-specific training parameters are listed in Table 5.
Table 4: Evaluation hyperparameters.
Dataset Maximum Length OTD prefix size OTD Cost $C$ T-mAP horizon T-mAP $\delta$ Transactions 32 5 1 7 (week) 2 MIMIC-IV 32 5 2 28 (days) 4 Retweet 32 10 15 180 (seconds) 30 Amazon 32 5 1 10 (unk. unit) 2 StackOverflow 32 10 1 10 (minutes) 2
Table 5: Training hyperparameters.
Dataset Num epochs Max Seq. Len. Label Emb. Size Hidden Size Head hiddens Transactions 30 1200 256 512 512 $\rightarrow$ 256 MIMIC-IV 30 64 16 64 64 Retweet 30 264 16 64 64 Amazon 60 94 32 64 64 StackOverflow 60 101 32 64 64
Figure 9: The dependency of metric values on the metric parameter.
Appendix F Labels Entropy Degradation
For simplicity, some datasets were omitted in the main part of the paper. In Figure 10, we show the dependency of label distribution entropy on step size for all datasets.
Figure 10: Entropy of label distributions as a function of the position in the generated sequence.
Appendix G The Optimal Sequence Length
For simplicity, some datasets were omitted from the main part of the paper. Figure 11 illustrates the relationship between label distribution entropy and step size across all datasets.
Figure 11: T-mAP dependency on the maximum length of the predicted sequence.

### v3

@@ C.1. Architecture
HoTPP incorporates best practices from extensible and reproducible ML pipelines. It leverages PyTorch Lightning (Falcon and The PyTorch Lightning team, 2019) as the core training library, ensuring reproducibility and portability across various computing environments. Additionally, HoTPP utilizes the Hydra configuration library (Yadan, 2019) to enhance extensibility. The overall architecture is illustrated in Fig. 7. HoTPP supports both discrete-time and continuous-time models as well as RNN and Transformer architectures. Implementing a new method requires only adding essential components, while the rest can be configured through Hydra and configuration files.
Figure 7. The architecture of the HoTPP library.
C.2. Metrics
@@
MTPP models are typically evaluated based on the accuracy of next-event predictions. Common metrics include mean absolute error (MAE) or root mean square error (RMSE) for time shifts and accuracy or error rate for label predictions. Some studies also assess test set likelihood as predicted by the model, but we do not include this measure as it is intractable for the IFTPP and Diffusion models. Previous works have advanced long-horizon evaluation using OTD (Mei et al., 2019; Xue et al., 2023). In addition to these metrics, we evaluate the novel T-mAP metric, which addresses the shortcomings of previous metrics, as discussed in Section 3.1.
C.3. Backbones
@@
We use three types of architectures in our experiments. The first is the GRU network (Cho et al., 2014), one of the top-performing neural architectures for event sequences (Babaev et al., 2022). We also implement the continuous-time LSTM (CT-LSTM) from the NHP method (Mei and Eisner, 2017). Since CT-LSTM requires a specialized loss function and increases training time, we use it exclusively with the NHP method, preferring GRU in other cases. An additional advantage of GRU is that its output is equal to its hidden state, simplifying the estimation of intermediate hidden states for autoregressive inference starting from the middle of a sequence. Lastly, we implement the AttNHP continuous-time transformer model (Yang et al., 2022).
C.4. Datasets
@@
Transactions [^3] https://huggingface.co/datasets/dllllb/age-group-prediction, Retweet [^4] https://huggingface.co/datasets/easytpp/retweet, Amazon [^5] https://huggingface.co/datasets/easytpp/amazon, and StackOverflow [^6] https://huggingface.co/datasets/easytpp/stackoverflow datasets were obtained from the HuggingFace repository. Transactions data was released in competition and came with a free license [^7] https://www.kaggle.com/competitions/clients-age-group/data. Retweet, Amazon, and StackOverflow come with an Apache 2.0 license. MIMIC-IV is subject to PhysioNet Credentialed Health Data License 1.5.0, which requires ethical use of this dataset. Because of a complex data structure, we implement a custom preprocessing pipeline for the MIMIC-IV dataset.
@@
C.5. Performance improvements
The HoTPP benchmark provides highly optimized training and inference procedures for the efficient evaluation of datasets containing up to tens of millions of events. First, we implement parallel RNN inference, which reuses computations when inference is initiated from multiple starting points within a sequence. Additionally, we optimize the code and apply PyTorch Script to the CT-LSTM model from NHP and the ODE model. CT-LSTM from EasyTPP and the ODE model from the original repository serve as baselines. The timing results are presented in Table 4. Experiments were conducted using a synthetic batch with a size of 64, a sequence length of 100, and an embedding dimension of 64. Evaluation was performed on a single Nvidia RTX 4060 GPU. The results show that HoTPP is 17 times faster at RNN autoregressive inference compared to simple prefix extension. HoTPP also accelerates CT-LSTM and ODE by 4 to 8 times. These optimizations significantly extend the applicability of the implemented methods to larger-scale datasets [^8] Experiments are documented in the “notebooks” folder of the HoTPP repository. Additionally, HoTPP offers a GPU implementation of the Hungarian algorithm, which also finds applications in computer vision.
Table 4. HoTPP computation speed improvements in terms of seconds per batch.
@@ Appendix D Training details
We trained each model using the Adam optimizer, which has a learning rate 0.001, and a scheduler that reduces the learning rate by 20% after every five epochs. Gradient clipping was applied, and the maximum L2-norm was set to 1.
We performed computations on NVIDIA V100 and A100 GPUs, with some smaller experiments conducted on an Nvidia RTX 4060. Each method was trained on a single GPU. The training time varied depending on the dataset and method, ranging from 5 minutes for RMTPP on the StackOverflow dataset to 40 minutes for the same method on the Transactions dataset and up to 15 hours for NHP on Transactions. Multi-seed evaluations took approximately five times longer to complete.
@@ Appendix E Hyperparameters
Dataset-specific training parameters are listed in Table 5.
Table 5. Training hyperparameters.
Dataset Num epochs Max Seq. Len. Label Emb. Size Hidden Size Head hiddens
Transactions 30 1200 256 512 512 $\rightarrow$ 256
MIMIC-IV 30 64 16 64 64
Retweet 30 264 16 64 64
Amazon 60 94 32 64 64
StackOverflow 60 101 32 64 64
Appendix F Datasets Analysis
HoTPP incorporates datasets from various domains, including finance, social networks, and healthcare. Below, we provide additional details highlighting the key differences among these datasets.
Ordering Sensitivity.
While most datasets—such as StackOverflow, Amazon, Retweet, and Transactions—exhibit a relatively balanced distribution of timestamps, we observed that in MIMIC-IV, the proportion of zero time steps exceeds 50%, as shown in Table 6. This results in a significant number of events sharing the same timestamp. Consequently, the actual order of events within the dataset may hold greater significance during evaluation than the precise prediction of timestamps. This characteristic can lead to instability in timestamp-based metrics, such as OTD and T-mAP, compared to index-based metrics like next-event quality, pairwise MAE and accuracy. The low next-event accuracy of methods based on NHP loss (NHP, AttNHP, and ODE) on MIMIC-IV can be attributed to their independent modeling of each event class, which results in a random ordering of events with identical timestamps.
Table 6. Time step percentiles.
Dataset 1% 5% 10% 50% 90% 95% 99%
Transactions 0.0 0.0 0.0 1.0 2.0 2.0 5.0
MIMIC-IV 0.0 0.0 0.0 0.0 6.9 63.8 768.0
Retweet 0.0 0.0 1.0 8.0 85.0 151.0 377.0
Amazon 0.01 0.01 0.01 0.73 0.79 0.79 0.80
StackOverflow 0.0003 0.01 0.05 0.51 2.16 2.97 5.08
Discrete Timestamps.
Some datasets feature continuous timestamps (e.g., StackOverflow, Amazon, MIMIC-IV), while others round timestamps to a specific precision (e.g., Retweet, Transactions). Modeling discrete timestamps presents a unique challenge, as density estimation methods like NHP can produce infinite PDF values. We introduce small Gaussian noise to discrete timestamps to address this issue, effectively smoothing the distribution. The degree of smoothing was carefully tuned for each dataset individually. The exact values are provided in the configuration files included with the source code.
Appendix G Qualitative Analysis of Predictions
Table 7 presents example predictions from various methods on the Transactions, MIMIC-IV, and Retweet datasets. The table focuses exclusively on predicted label sequences. It can be seen that all methods, except HYPRO, exhibit issues with constant or repetitive outputs. We believe this behavior stems from a bias in the predictions toward the most frequent labels, especially in scenarios with high uncertainty. For the MIMIC-IV dataset, the prediction patterns differ significantly, as most events follow a predefined sequence, such as admission, a standard set of laboratory tests, and diagnosis. In this case, the uncertainty is lower, enabling the methods to generate sequences with more diverse event types. Addressing the mentioned limitation in future research could lead to the development of methods capable of producing more varied and realistic sequences in high-uncertainty scenarios.
Table 7. Example predictions (labels only).
Method Seq. ID Transactions MIMIC-IV Retweet
IFTPP 0 3, 1, 3, 1, 3, 1, 3, 3, 3 10, 27, 23, 22, 1, 27, 3, 28, 26, 25, 23, 30 1, 1
1 6, 3, 6, 6, 6, 6, 6 11, 10 0, 0, 0, 0, 0, 0, 0, 0, 0
2 3, 1, 3, 1, 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
RMTPP 0 3, 1, 3, 1, 3, 1 10, 27, 23, 22, 1, 30, 27, 5, 8, 9, 16, 15 1, 1, 1, 1
1 6, 6, 6, 6, 6, 6 11, 10, 1 0, 0, 0, 0, 0, 0, 0, 0
2 3, 1, 3, 1, 3 2, 7, 14, 12, 4, 6, 11, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
NHP 0 1, 1, 1, 1, 1, 3, 1, 1, 1, 1 6 1, 1, 1, 1
1 1, 1, 6, 6, 6, 6, 6 6, 6, 6 1, 1, 1, 1, 1, 1, 1
2 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 6, 6 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
ODE 0 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1 10, 1, 3, 5, 2, 4, 6, 1 1, 1, 1, 1
1 6, 6, 6, 6, 6 1, 3, 5, 2, 4, 6, 1, 3, 5, 2, 4, 6 0, 0, 0, 0, 0, 0, 0
2 1, 1, 1, 1, 1, 1, 1 2, 4, 6, 1, 3, 5, 2, 4, 10 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1
HYPRO 0 3, 1, 16, 3, 12, 1 23, 27, 22, 1, 28, 26, 25, 23 0, 0, 1, 0
1 1, 32, 6, 6, 1, 6 1, 28, 25, 23, 3, 26, 22, 5 0, 0, 0, 1, 1, 0, 1
2 3, 5, 1, 3, 5 15, 13, 2, 7, 4, 1, 3, 5 0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0
Diffusion 0 1, 3, 1, 1, 1, 3, 18, 1 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0
1 6, 6, 6, 6, 6, 6, 6 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 0, 0, 1, 1, 1, 1, 0, 1, 0
2 8, 1, 3, 8, 3, 8, 1, 3, 8, 1, 1, 8 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0 1, 1, 1, 1, 1, 0, 1, 1, 1, 1, 1, 1
Appendix H T-mAP for Highly Irregular Sequences
This section analyzes a toy dataset containing highly irregular event sequences. This dataset includes a single label, which aims to predict timestamps. Most time intervals in the dataset are zero, with only 5% of intervals equal to one. We compare three simple baselines:
  • ZeroStep, which predicts events with timestamps identical to the last observed event (zero intervals);
  • UnitStep, which predicts events with a unit time step (the largest time step in the dataset);
  • MeanStep predicts events using the average time step computed from historical data.
Evaluation results are shown in Fig. 8. The results indicate that the MAE and OTD metrics assign the lowest error to the ZeroStep baseline, which simply repeats the last event without accounting for the dataset’s irregularity. In contrast, T-mAP identifies the MeanStep baseline as the most effective, as it is the only method that analyzes historical data and incorporates timestamp statistics (mean interval) into its predictions.
These findings suggest that T-mAP is a more appropriate metric for assessing the ability of methods to predict irregular events.
Figure 8. Comparison of simple baselines on the Toy dataset with highly irregular time intervals. MAE and OTD metrics represent error values, while T-mAP measures model quality.
Appendix I Long-tail Prediction
In this section, we assess the capability of various evaluation metrics to capture long-tail prediction quality, specifically the ability of models to predict rare classes. Unlike OTD, T-mAP evaluates each class independently, allowing for different aggregation strategies. The standard T-mAP computes a macro average, where the quality for each event class contributes equally to the final score. Additionally, the HoTPP benchmark includes a weighted variant of T-mAP, where the weights are proportional to class frequencies. Fig. 9 compares the performance of IFTPP and RMTPP on the Transactions dataset, which includes 203 classes. The results show that all metrics, except macro T-mAP, remain unaffected as the dataset size increases beyond 60 classes. In contrast, macro T-mAP effectively evaluates the ability of models to predict across all available classes, highlighting its suitability for long-tail prediction tasks.
Figure 9. Comparison of various metrics on the Transactions dataset across subsets with fewer event classes.